# AdaSemiCD: An Adaptive Semi-Supervised Change Detection Method Based on Pseudo-Label Evaluation

### v1

@@ Abstract
Change Detection (CD) is an essential field in remote sensing, with a primary focus on identifying areas of change in bi-temporal image pairs captured at varying intervals of the same region by a satellite. The data annotation process for the CD task is both time-consuming and labor-intensive. To make better use of the scarce labeled data and abundant unlabeled data, we present an adaptive dynamic semi-supervised learning method, AdaSemiCD, to improve the use of pseudo-labels and optimize the training process. Initially, due to the extreme class imbalance inherent in CD, the model is more inclined to focus on the background class, and it is easy to confuse the boundary of the target object. Considering these two points, we develop a measurable evaluation metric for pseudo-labels that enhances the representation of information entropy by class rebalancing and amplification of confusing areas to give a larger weight to prospects change objects. Subsequently, to enhance the reliability of sample-wise pseudo-labels, we introduce the AdaFusion module, which is capable of dynamically identifying the most uncertain region and substituting it with more trustworthy content. Lastly, to ensure better training stability, we introduce the AdaEMA module, which updates the teacher model using only batches of trusted samples. Experimental results from LEVIR-CD, WHU-CD, and CDD datasets validate the efficacy and universality of our proposed adaptive training framework.
@@
Change Detection, Semi-supervised Learning, Remote Sensing, Mean-Teacher
@@ I Introduction
Change detection (CD) has emerged as a significant research focus within the field of remote sensing in recent years. Its objective is to identify regions of interest that have experienced alterations in bi-temporal image pairs captured at varying times of the same geographical area using satellite technology. This method plays a crucial role in remote sensing data analysis and is of particular importance in various civilian sectors such as urban planning [1, 2], rural land management [3, 4], and disaster assessment [5, 6].
Figure 1: Comparison between SOTA semi-supervised methods and our AdaSemiCD on the CDD Dataset.
@@
Given that the process of accurately annotating masks for change detection tasks is notably labor intensive, direct application of traditional supervised learning approaches, such as CNN [7, 3, 8, 9, 4] and Transformers [10, 11, 12], to a limited set of labeled data often results in limited performance. In response to these challenges, researchers have investigated a range of approaches such as Self-supervised learning(SSL)[13, 14],Unsupervised Change Detection (USCD) [15, 16, 17], Weakly Supervised Change Detection (WSCD) [18, 19, 20], and sample generation strategies [21, 22, 23, 24]. Although WSCD is cost-efficient, it is reliant on incomplete or inaccurate labels, which can introduce substantial errors and unpredictably noisy data. USCD, on the other hand, does not require labeled data and utilizes the intrinsic patterns present in the data, but it often faces challenges when tackling specific tasks like classification or detection. Sample generation strategies, which include data augmentation [23], generative adversarial networks (GANs) [22], and diffusion models [24], frequently necessitate the simulation or synthesis of additional data. However, when dealing with limited available samples, these methods may encounter constraints due to insufficient diversity in the generated data, which can diminish the model’s ability to generalize. As a result, semi-supervised change detection (SSCD) [25, 26, 27] emerges as a potentially more effective solution. The paradigm of semi-supervised learning [28, 29, 30] is to enhance CD performance by leveraging the limited available labels and the large volume of unlabeled samples. Usually, researchers generate pseudo-labels for the unlabeled data to act as guidance during training. These pseudo-labels are often temporary predictions with higher probabilities. The most prevalent approach is the Mean-Teacher [31] framework, which employ a teacher model to generate pseudo-labels that serve as guidance for the student model during the training process. The teacher model is subsequently updated using the exponential moving average (EMA) [32] of the student model. By training on a mix of limited data with actual labels and abundant data with pseudo-labels, the student model can learn more significant features, leading to noticeable enhancements in performance.
@@
While these methods produce acceptable outcomes, significant problems persist: the model indiscriminately treats all samples, irrespective of their quality, and the training process lacks flexibility. Firstly, it’s evident that unlabeled samples may not always function as efficient ‘teachers’. Models frequently encounter difficulties in generating reliable high-quality pseudo-labels for intricate samples, which in turn introduces extra noise that can misguide the model’s training.
@@
Subsequently, the EMA updating process does not take into account the quality of samples. Given that training batches can be biased or contain noise, dynamically determining the training update timing could contribute to the stability of the training process. These factors underscore the need for a more precise supervisory approach, failing which it could negatively impact the model’s training. In this study, we introduce an adaptive dynamic learning strategy, AdaSemiCD, designed to improve the accuracy of pseudo-labels and streamline the training process. Our framework incorporates the traditional semi-supervised training approach, augmented by two innovative functional modules, AdaFusion and AdaEMA. Initially, we utilize AdaFusion to suppress noise at the individual sample level, thereby enhancing the accuracy of pseudo-labels. Contrary to previous methods like Augseg [33] or CutMix [34] that relied on entirely random fusion regions, our AdaFusion technique proactively identifies the most uncertain region and substitutes it with reliable content from either labeled datasets or unlabeled datasets of superior quality. Following this, we dynamically adjust the rate of parameter updates in the teacher-student model via AdaEMA to ensure improved stability. Although the traditional EMA effectively mitigates fluctuations in model parameters, thereby boosting stability, it persists in uniformly updating after each training iteration, overlooking the model’s varying learning outcomes across different iterations when handling a range of training samples. If unlabeled samples contain an abundance of erroneous information, it can misdirect the model’s training. Therefore, our AdaEMA introduces an adaptive selection process for model-level parameter updates, enabling the model to fully integrate superior parameters.
In summary, the main contributions of this paper are as follows.
@@
  • Experimental results on publicly available datasets LEVIR-CD, WHU-CD, and CDD demonstrate the effectiveness of our method.
@@ II-A Semi-supervised learning
The technique of semi-supervised learning (SSL) applies supervised learning to a small subset of labeled data and unsupervised learning to a larger pool of unlabeled data. SSL is mainly categorized into three strategies: consistent regularization (CR), self-training, and holistic methods. The last strategy combines the first two strategies within an SSL framework.
CR techniques are grounded in the concept of perturbed consistency, which utilizes the coherence between the model’s output after varying degrees of perturbed input data as a training constraint. The three consistent regularization frameworks, in ascending order, including: the ${\Pi}$-model [35], the Temporal-ensembling model [35], and the Mean-teacher (MT) model [32]. The ${\Pi}$-model’s double-branch network shares the weight; the Temporal-ensembling model amalgamates all the outputs in the time series, with each image’s pseudo-labels being the EMA of the previously generated results; the MT model carries out this smoothing operation at the model parameters level. This model has found application in subsequent semi-supervised research across various domains, such as Active-Teacher for semi-supervised object detection [36], [37, 38, 33, 39] for semi-supervised general semantic segmentation,  [40] for image classification, and  [41][42] for semi-supervised medical image segmentation. There has also been explored in the area of perturbation design, with [43] and [25] examining the image-level perturbation and feature-level perturbation of CR respectively.
In the realm of self-training methods, the authenticity of pseudo-labels is of paramount importance. This has led to extensive research into the effective selection of high-quality pseudo-labels for supervised learning. [44] employs a set probability threshold as a selection standard. ST++ [45] has developed a multi-tier self-training structure, where labels of high confidence are used for self-training repeatedly until all unlabeled samples have been utilized. [46] uses a constant entropy value as the filtering limit.
More often than not, CR and self-training are not used independently, but rather in conjunction, forming a comprehensive approach to semi-supervised learning. This is exemplified in [38, 33, 39, 40] and [45], as partially discussed in earlier sections.
@@ II-B Semi-supervised Change Detection
Since annotating a large number of images for CD is time-consuming, recent methods mainly focus on the SSCD. In the realm of CR techniques, the incorporation of the mean teacher model in CD was first introduced by Bousias et al. [47]. However, the initial outcomes did not show considerable potential, as this SSCD method fell short when compared to a benchmark that exclusively used a restricted quantity of labeled data for entirely supervised learning. Despite increasing labeled data, this disparity continues to expand. Using this as a basis, Mao et al.[48] implemented minor and major improvements to the inputs of the teacher and student models, respectively. Furthermore, they formulated an extra teacher-virtual adversararial training component to further reduce the harmful effects of the pseudo label noise.
@@
Additionally, other semi-supervised methods employ either a single model or a two-branch model with shared weights. Such as Sun et al. [49] introduce a siamese network. They incorporated additional self-training based on pseudo-labels, employing threshold filtering to eliminate low-quality pseudo-labels. The rationale behind this filtering lies in the potential noise introduced by pseudo-labels with low confidence, which could adversely affect SSL training. Hafner et al. [50] propose a dual-task SSCD framework that combines building segmentation and change detection, two closely related downstream tasks. They devised a novel consistency constraint between the two change detection masks produced by the Siamese segmentation network and the change detection network. Bandara et al. [25] explored feature-based perturbations of the regularization term, applying various data perturbations at the feature level to expand the distribution space of consistency constraints. This approach fully leverages the information embedded in unlabeled samples, and in recent work, Zhang et al. [27] imposed two constraints of class consistency and feature consistency on unlabeled datasets. By aligning the feature representations of unlabeled samples on varying and invariant classes, the model could learn from a feature space that is closer to the real distribution, this contributed to the renewal of the best performance record at that time.
Other methods predominantly utilize Generative Adversarial Network (GAN), which were initially proposed by Ian J. Goodfellow in 2014[51]. Some methods use GAN to learn the feature distribution space close to the real labeled data[26, 52, 53, 54]; The other part uses GAN to generate data samples[55, 56]; The more noteworthy research in the recent work is [18] proposed a new paradigm for change detection, unifying unsupervised, weakly supervised, regionally supervised, and fully supervised change detection into an end-to-end framework. While these endevavors have reported great success in SSCD, the highly unstable training of GAN makes hyperparameter adjustment become s challenging. Additionally, the issue of gradient disapperance frequently arises during the training phase. Morever, the strong discriminatory capaality of the discriminator can lead to imbalanced performance between the generator and discriminator of the GAN if no additional training techniques are implementrd. Consquently, achieving the idealized optimal scenario is challenging, making the pritical application of this method somewhat intricate. So we continue the semi-supervised framework of consistent and self-teinging.
@@ III METHODOLOGY
Fig. 2 provides a comprehensive summary of our AdaSemiCD framework, aiming to enhance the SSCD performance by utilizing the scarce labels and vast quantity of unlabeled samples. We commence with a broad introduction of the framework, followed by an in-depth explanation of the uncertainty map used to assess our pseudo-labels, and finally present the specifics of AdaFusion and AdaEMA.
Figure 2: The overview of the proposed AdaSemiCD framework, and details of two novel modules. Our approach is based on the common teacher-student training pipeline for SSCD, and we suggest utilizing the AdaFusion module to produce samples with increased reliability and the AdaEMA module to enhance the efficiency of EMA update times.
@@ III-A Overview of the AdaSemiCD Framework
The task of semi-supervised change detection can be described as follows: Given a labeled dataset ${D}_{l}=\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}_{i=1}^{m}$ and an unlabeled dataset ${D}_{u}=\left\{x_{a,j}^{u},x_{b,j}^{u}\right\}_{j=1}^{n}$, where $\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}$ represents the i-th pair of labeled image pairs and their ground truth, and $\left\{\smash{x_{a,i}^{u},x_{b,i}^{u}}\right\}$ represents the i-th pair of unlabeled image pairs, subscripts a and b are used to identify the images before and after phases respectively, and the number of samples with and without labeled datasets is n and m, $n\gg m$, the model $M$ is expected to not only extract meaningful information from ${D}_{l}$ but also capture a wider range of features by utilizing numerous unlabeled training samples in ${D}_{u}$ to improve the model’s generalization capabilities.
Model architecture: In this study, we employ the widely-used teacher-student framework for SSCD tasks. The network is composed of two components, a student model $M_{stu}$ and a teacher model $M_{tea}$ that shares the same structure. The student model is trained to extract significant features from a small number of labeled samples and a large volume of unlabeled samples, with the aid of optimization via gradient descent methods. Conversely, the teacher model $M_{tea}$ generates pseudo-labels to guide the student in assimilating unlabeled data, and the teacher model is refreshed using EMA methods.
Usually, samples undergo either weak $A_{w}(\cdot)$ or strong $A_{s}(\cdot)$ augmentation processes prior to being input into the network to maintain high-quality generalization performance.
Objectives: The objective is to minimize the supervised loss $\mathcal{L}_{s}$ on $D_{l}$ while also ensuring consistency on disturbed $D_{u}$ with a minimal $\mathcal{L}_{u}$. During training, samples are fed into the network in randomly shuffled batches, $\mathcal{B}_{l}$ and $\mathcal{B}_{u}$.
The loss $\mathcal{L}_{s}$ for labeled samples within a batch $B_{l}$ is calculated as the cross entropy (CE) of ground truth $y_{i}^{l}$ and it’s prediction $p_{i}^{l}$:
@@
where $|\mathcal{B}_{l}|$ represents the mini-batch’s size, ${p}_{i}^{l}$ represents the CD result of the $i$-th pair images.
The loss $\mathcal{L}_{u}$ for unlabeled samples within a batch $B_{u}$ is quite similar. Here we use the pseudo-labels from $M_{tea}$ as supervision for the prediction from $M_{stu}$. And $\mathcal{L}_{u}$ is calculated as:
@@
where $|\mathcal{B}_{u}|$ represents the size of unlabeled image mini-batch, ${p}_{w,j}^{u}=M_{tea}^{\theta^{\prime}}(w_{a,j}^{u},w_{b,j}^{u})$ represents the detection result of the teacher model on the change of the i-th pair of unlabeled images weakly augmented, and ${p}_{s,j}^{u}={M_{stu}^{\theta}(s_{a,j}^{u},s_{b,j}^{u})}$ represents the detection result of the teacher model on the change of the i-th pair of unlabeled images by further strongly augmented.
To sum up, the total loss for the training procedure of AdaSemiCD is:
$\mathcal{L}=\mathcal{L}_{s}+\lambda(\cdot)\mathcal{L}_{u}$ (3)
Previous studies[40, 41] have employed a constant value for $\lambda(\cdot)$. However, we argue that this approach is not suitable for single-stage semi-supervised learning algorithms. During the initial training phases, the pseudo-labels generated by our model for unlabeled samples are highly unreliable; excessive reliance on unsupervised training at this stage can introduce significant noise. Conversely, as training progresses into the middle and later stages, the model increasingly learns from the limited labeled data, leading to an improvement in the quality of the pseudo-labels. This justifies a reduction in the proportion of supervised training relative to unsupervised training. To mitigate overfitting and enhance the feature space, it is essential to implement a controlled process to dynamically adjust the contribution of these two components in the loss function. As the pseudo-labels in the starting stage are quite unstable, we use a ramp-up process to control the training pace of the unlabeled part. $\lambda(\cdot)$ denotes the time varing weight of unsupervised loss, which is a function changing with training step and used to dynamically adjust the proportion of supervised and unsupervised training in different stages.
@@
where $w_{max}$ and $\phi$ are hyperparameters, $w_{max}$ representing the maximum weight of unsupervised loss, $\phi$ is used to control the severity of the ramp-up; $iter_{cur}$ represents the current iteration cycles; $iter_{max}$ is the total number of ramp-up cycles, is obtained by multiplying $\gamma$ and the total number of training iterations, and $0\textless\gamma\textless 1.0$; After the ramp-up process, the weight of unsupervised losses remains $w_{max}$ and does not change. In the early stage of training, this weight is relatively low, and unsupervised training plays a negligible role, but in the middle and later stages of training, this weight gradually increases and the unsupervised loss after weighting exceeds the supervised loss, so unsupervised training accounts for the main part.
Training strategy: The student network’s parameters $\theta$ are adjusted by Stochastic Gradient Descent(SGD) technique to reduce the overall loss $\mathcal{L}$, while the teacher network’s parameters $\theta^{\prime}$ are updated by the exponential moving average of the student model’s parameters $\theta$ over the time series, as depicted in Eq. 5. The hyperparameter $\beta$ acts as a momentum parameter. The larger the value of $\beta$, the broader the moving metric average window becomes. Typically, $\beta$ is chosen to be a number near 1.0, for instance, 0.996 in this study.
$\theta^{\prime}=\beta\theta^{\prime}+(1-\beta)\theta$ (5)
Proposed modules: The essence of semi-supervised learning lies in the quality of pseudo-labels. Nevertheless, it’s apparent that the previously mentioned procedure does not take into account the diverse effects of individual samples on training. This paper concentrates on two elements that are directly related to the generation of pseudo-labels: the initial pair of unlabeled images, and the efficiency of the pseudo-label-generation network, also known as the teacher model, in identifying changes. To offer more dependable supervisory data to unlabeled information and reduce training uncertainty, we’ve designed an adaptive training strategy to tackle these two key issues. Our initial proposal is to develop a metric that can measure the uncertainty of pseudo-labels, serving as the basis for adaptive modifications. Following that, we suggest making adaptive changes at the image level to the unlabeled training samples and suitably integrating reliable contents. Furthermore, it includes applying adaptive and selective EMA updates to the teacher network during the training phase to limit variations, thus producing more consistent and superior quality pseudo-labels. These efforts are elaborated in the subsequent sections.

### v2

# AdaSemiCD: An Adaptive Semi-supervised Change Detection Method Based on Pseudo Label Evaluation

@@ Abstract
Change detection(CD) is an essential field in remote sensing, with a primary focus on identifying areas of change in bitemporal image pairs captured at varying intervals of the same region. The data annotation process for CD tasks is both time-consuming and labor-intensive. To better utilize the scarce labeled data and abundant unlabeled data, we introduce an adaptive semi-supervised learning method, AdaSemiCD, to improve pseudo-label usage and optimize the training process. Initially, due to the extreme class imbalance inherent in CD, the model is more inclined to focus on the background class, and it is easy to confuse the boundary of the target object. Considering these two points, we develop a measurable evaluation metric for pseudo-labels that enhances the representation of information entropy by class rebalancing and amplification of ambiguous areas, assigning greater weights to prospective change objects. Subsequently, to enhance the reliability of samplewise pseudo-labels, we introduce the AdaFusion module, to dynamically identifying the most uncertain region and substituting it with more trustworthy content. Lastly, to ensure better training stability, we introduce the AdaEMA module, which updates the teacher model using only batches of trusted samples. Experimental results on ten public CD datasets validate the efficacy and generalizability of our proposed adaptive training framework.
@@
Pseudo label, Semi-supervised Learning, Change Detection, Mean Teacher, Adaptive Learning
@@ I Introduction
Change detection (CD) has emerged as a significant research focus within the field of remote sensing. Its objective is to identify regions of interest that have experienced alterations in bi-temporal image pairs captured at varying times of the same geographical area. This method plays a crucial role in remote sensing data analysis and is particularly important in various civilian sectors, such as urban planning [1, 2], rural land management [3, 4], and disaster assessment [5, 6].
@@
Given that the process of accurately annotating masks for change detection tasks is notably labor-intensive, direct application of traditional supervised learning approaches, such as convolutional neural networks (CNN) [7, 8] and Transformers [9, 10], to a limited set of labeled data often results in limited performance. In response to these challenges, researchers have explored a range of approaches such as weakly supervised change detection (WSCD) [11, 12], unsupervised change detection (USCD) [13, 14], and semi-supervised change detection (SSCD) [15, 16]. Although WSCD is cost-efficient, it relies on incomplete or inaccurate labels, which can introduce significant errors and unpredictable noise. USCD, on the other hand, does not require labeled data and leverages the intrinsic patterns present in the data; however, it often faces challenges when tackling specific tasks like classification or detection. Some methods would adopt sample generation strategies [17, 18, 19, 20], which include data augmentation [19], generative adversarial networks (GAN) [18], and diffusion models [20], frequently necessitate the simulation or synthesis of additional data. However, when dealing with limited available samples, these methods may encounter constraints due to insufficient diversity in the generated data, which can diminish the model’s ability to generalize. As a result, SSCD [21, 22, 23] emerges as a potentially more effective solution. The paradigm of semi-supervised learning (SSL) [24, 25, 26] aims to enhance CD performance by leveraging the limited available labeled data and the large volume of unlabeled samples. Typically, researchers generate pseudo-labels for the unlabeled data to act as guidance during training. These pseudo-labels are often temporary predictions with higher probabilities. The most widely used approach is the mean-teacher (MT) [27] framework, which employs a teacher model to generate pseudo-labels that serve as guidance for the student model during the training process. The teacher model is subsequently updated using the exponential moving average (EMA) [28] of the student model. The student model benefits from training on a mix of limited labeled data along with ample pseudo-labeled data, enabling it to identify more important features and resulting in marked enhancements in performance.
@@
While these methods produce acceptable outcomes, significant problems persist: the model indiscriminately treats all samples, irrespective of their quality, and the training process lacks flexibility. Firstly, it is evident that unlabeled samples may not always function as efficient ‘teachers’. Models frequently encounter difficulties in generating reliable high-quality pseudo-labels for intricate samples, which in turn introduces extra noise that can mislead the model’s training. Subsequently, the EMA updating process does not take into account the quality of samples. Given that training batches may be biased or contain noise [29], dynamically determining the timing of training updates could contribute to the stability of the training process. These factors underscore the need for a more precise supervisory approach, failing which it could negatively impact the model’s training.
@@
In this study, we introduce an adaptive learning strategy, AdaSemiCD, designed to improve the accuracy of pseudo-labels and streamline the training process. Our framework builds upon the traditional semi-supervised training approach, augmented by two innovative functional modules, AdaFusion and AdaEMA. Initially, AdaFusion is employed to suppress noise at the individual sample level, thereby enhancing the accuracy of pseudo-labels. Contrary to previous methods like Augseg [30] or CutMix [31] that relied on entirely random fusion regions, our AdaFusion technique proactively identifies the most uncertain region and substitutes them with reliable content from either labeled datasets or unlabeled datasets of superior quality. Following this, we dynamically adjust the rate of parameter updates in the teacher-student model via AdaEMA to ensure improved stability. Although the traditional EMA effectively mitigates fluctuations in model parameters, thereby boosting stability, it persists in uniformly updating after each training iteration, neglecting the model’s varying learning outcomes across different iterations when handling a range of training samples. If unlabeled samples contain excessive erroneous information, it can misdirect the model’s training. Therefore, our AdaEMA introduces an adaptive selection process for model-level parameter updates, allowing the model to fully integrate superior parameters.
The main contributions of this paper are as follows:
@@
  • Experimental results on ten publicly available datasets demonstrate the effectiveness of our method.
@@ II-A Semi-supervised Learning
Semi-supervised learning involves applying supervised learning on a limited amount of labeled data while employing unsupervised learning on a vast set of unlabeled data. SSL is typically divided into three strategies: consistent regularization (CR), self-training, and holistic methods, with the latter integrating the first two strategies within an SSL framework.
CR techniques are grounded in the concept of perturbed consistency, which utilizes the coherence between the model’s output after varying degrees of perturbed input data as a training constraint. The three typical consistent regularization frameworks consist of the ${\Pi}$-model [32], the Temporal-ensembling model [32], and the mean-teacher model [28]. The ${\Pi}$-model’s double-branch network shares the weight; the Temporal-ensembling model amalgamates all the outputs in the time series, with each image’s pseudo-labels being the EMA of the previously generated results; the MT model carries out this smoothing operation at the model parameters level. This model has found application in subsequent semi-supervised research across various domains, such as Active-Teacher for semi-supervised object detection [33], [34, 35, 36] for semi-supervised general semantic segmentation, [37] for image classification, and [38, 39] for semi-supervised medical image segmentation. There has also been explored in the area of perturbation design, with [40] and [21] examining the image-level and feature-level perturbation of CR respectively.
In the realm of self-training methods, the authenticity of pseudo-labels is of paramount importance. This has led to extensive research into the effective selection of high-quality pseudo-labels for supervised learning. Feng et al.[41] proposed a method that incrementally adds labeled instances, reduces class bias via Synthetic minority over-sampling technique, and adaptively selects the optimal number of instances to enhance classifier performance. [42] employs a set probability threshold as a selection standard. ST++ [43] has developed a multi-tier self-training structure, where labels of high confidence are used for self-training repeatedly until all unlabeled samples have been utilized. [44] uses a constant entropy value as the filtering limit.
CR and self-training are commonly employed together rather than in isolation, creating a holistic strategy for semi-supervised learning. This is illustrated in [34, 30, 36, 37] and [43], as partially described in previous sections.
@@ II-B Semi-supervised Change Detection
Since annotating a large number of images for CD is time-consuming, recent methods mainly focus on the SSCD. In the realm of CR techniques, the incorporation of the mean-teacher model in CD was first introduced by Bousias et al. [45]. However, the initial outcomes did not show considerable potential, as this SSCD method fell short when compared to a benchmark that exclusively used a restricted quantity of labeled data for entirely supervised learning. Despite increasing labeled data, this disparity continues to expand. Using this as a basis, Mao et al.[46] implemented minor and major improvements to the inputs of the teacher and student models, respectively. Furthermore, they formulated an extra teacher-virtual adversararial training component to further reduce the harmful effects of the pseudo label noise.
@@
Additionally, other semi-supervised methods employ either a single model or a two-branch model with shared weights. Such as Sun et al. [47] introduce a siamese network. They incorporated additional self-training based on pseudo-labels, employing threshold filtering to eliminate low-quality pseudo-labels. The rationale behind this filtering lies in the potential noise introduced by pseudo-labels with low confidence, which could adversely affect SSL training. Hafner et al. [48] propose a dual-task SSCD framework that combines building segmentation and change detection, two closely related downstream tasks. They devised a novel consistency constraint between the two change detection masks produced by the siamese segmentation network and the CD network. Bandara et al. [21] explored feature-based perturbations of the regularization term, applying various data perturbations at the feature level to expand the distribution space of consistency constraints. This approach fully leverages the information embedded in unlabeled samples. In recent work, Zhang et al. [23] imposed two constraints of class consistency and feature consistency on unlabeled datasets. By aligning the feature representations of unlabeled samples on varying and invariant classes, the model could learn from a feature space that is closer to the real distribution, this contributed to the renewal of the best performance record at that time.
Some approaches primarily employ generative adversarial networks to produce data samples and understand feature distributions that approximate real labeled data [22, 49, 50]. Although these efforts have shown notable achievements in SSCD, the notoriously erratic nature of GAN training presents difficulties with hyperparameter tuning. Furthermore, the occurrence of gradient vanishing is a common challenge during training. Moreover, the discriminator’s robust ability to differentiate can cause an imbalance between the generator and discriminator’s performance within the GAN unless additional training strategies are applied. As a result, reaching an ideal balance is demanding, complicating the practical application of this method. Therefore, we continue the semi-supervised framework leveraging consistency and self-training techniques.
@@ III METHODOLOGY
Fig. 1 provides a comprehensive summary of our AdaSemiCD framework, aiming to enhance the SSCD performance by leveraging scarce labeled data and a vast quantity of unlabeled samples. We commence with a general introduction of the framework, followed by an detailed explanation of the uncertainty map used to assess our pseudo-labels, and finally present the specifics of AdaFusion and AdaEMA.
Figure 1: The overview of the proposed AdaSemiCD framework and details of two adaptive modules. Our approach is based on the common MT training pipeline for SSCD, and we suggest utilizing the AdaFusion module to produce samples with increased reliability and the AdaEMA module to enhance the efficiency of EMA update.
@@ III-A Overview of the AdaSemiCD Framework
The process of semi-supervised change detection is broadly outlined as follows: We are given a labeled dataset, denoted as ${D}_{l}=\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}_{i=1}^{m}$, and an unlabeled dataset ${D}_{u}=\left\{x_{a,j}^{u},x_{b,j}^{u}\right\}_{j=1}^{n}$. Here, $\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}$ illustrates the $i$-th pair of labeled images alongside their corresponding true labels, while $\left\{\smash{x_{a,j}^{u},x_{b,j}^{u}}\right\}$ refers to the $j$-th pair of unlabeled images. The subscript $a$ signifies the image from the ‘Pre’-event period, and $b$ signifies the ‘Post’-event period. Importantly, the quantity of samples with labels and those without are $m$ and $n$, with $n$ considerably larger than $m$. The aim is for model $M$ to not only derive key insights from ${D}_{l}$ but also to enhance its feature extraction capability using the extensive collection of unlabeled samples in ${D}_{u}$, thus boosting the model’s generalization potential. Ordinarily, samples are subject to either weak augmentation $A_{w}(\cdot)$ or strong augmentation $A_{s}(\cdot)$ prior to being passed into the network to guarantee superior generalization performance.
Model architecture: In this study, we employ the widely used MT framework for SSCD tasks. Two integral parts make up the network: a student model, denoted as $M_{stu}$, and a teacher model, represented as $M_{tea}$. Both components possess an identical architecture. The student model is trained to extract significant features from a small number of labeled samples and a large volume of unlabeled samples, with the aid of optimization via gradient descent methods. Conversely, the teacher model $M_{tea}$ generates pseudo-labels to guide the student in assimilating unlabeled data, and it is updated using the EMA method.
Objectives: The objective is to minimize the supervised loss $\mathcal{L}_{s}$ on $D_{l}$ while ensuring consistency on the disturbed $D_{u}$ with a minimal $\mathcal{L}_{u}$. During training, samples are fed into the network in randomly shuffled batches, $\mathcal{B}_{l}$ and $\mathcal{B}_{u}$.
The loss $\mathcal{L}_{s}$ for labeled samples within a batch $\mathcal{B}_{l}$ is calculated as the cross entropy(CE) of ground truth $y_{i}^{l}$ and its prediction $p_{i}^{l}$:
@@
where $|\mathcal{B}_{l}|$ represents the mini-batch size, ${p}_{i}^{l}$ represents the change detection result for the $i$-th pair of images.
The loss $\mathcal{L}_{u}$ for unlabeled samples within a batch $B_{u}$ is quite similar. Here, we use the pseudo-labels from $M_{tea}$ to supervise the predictions from $M_{stu}$. The loss $\mathcal{L}_{u}$ is calculated as follows:
@@
where $|\mathcal{B}_{u}|$ represents the size of the unlabeled image mini-batch. ${p}_{w,j}^{u}=M_{tea}^{\theta}(w_{a,j}^{u},w_{b,j}^{u})$ and ${p}_{s,j}^{u}=M_{stu}^{\theta}(s_{a,j}^{u},s_{b,j}^{u})$ denote the change detection outcomes from the teacher model for the $j$-th pair of weak and strong augmentations of unlabeled images, respectively.
To summarize, the overall loss associated with the AdaSemiCD training process is defined as:
$\mathcal{L}=\mathcal{L}_{s}+\lambda(\cdot)\mathcal{L}_{u}.$ (3)
$\lambda$ is the weight of unsupervised loss, which is typically set to a constant value [37, 38]. For SSCD, we contend that using a constant $\lambda$ could potentially disrupt the training procedure. In the early stages of training, the pseudo-labels for unlabeled data tend to be highly unreliable, and excessive dependence on unsupervised training at this point can inject substantial noise. Conversely, as training progresses into the intermediate and final phases, the model enhances its ability to generate high-quality pseudo labels, diminishing the importance of the limited labeled dataset. This shift warrants a gradual decrease in the emphasis on supervised training compared to unsupervised training over time. To prevent overfitting and refine the feature space, it is crucial to have a systematic approach that dynamically modifies the balance between these two elements in the loss function. We employ a ramp-up approach, defining $\lambda(\cdot)$ as a function that adapts through the course of training, thereby dynamically modulating the balance between supervised and unsupervised training throughout various phases.
@@
$iter_{max}=\gamma\times iter_{t}otal$ (5)
Here, $w_{max}$ represents the maximum weight value of the unsupervised loss, and $\phi$ controls the severity of the ramp-up. $iter_{cur}$ represents the current iteration cycles; $iter_{max}$ is the total number of ramp-up cycles, calculated by multiplying $\gamma$ (where $10<\gamma<$ 1) by the total number of training iterations, as shown in 5. After the ramp-up process, the weight of the unsupervised loss stabilizes at $w_{max}$ and no longer changes. In the early stages of training, this weight is relatively low, so unsupervised training plays a negligible role, but in the middle and later stages of training, this weight gradually increases and the unsupervised loss after weighting exceeds the supervised loss, making unsupervised training the dominant factor.
Training strategy: To reduce the total loss $\mathcal{L}$, the parameters of the student network, denoted as $\theta_{stu}$, are refined via Stochastic Gradient Descent (SGD). Concurrently, the teacher network updates its parameters $\theta_{tea}$ using an exponential moving average calculated from the student network’s parameters $\theta_{stu}$ over a time sequence:
$\theta_{tea}=\beta\theta_{tea}+(1-\beta)\theta_{stu}$ (6)
The hyperparameter $\beta$ acts as a momentum factor, where a higher $\beta$ value leads to a broader moving average window. Generally, $\beta$ is selected to be near 1.0; in this study, for instance, $\beta$ is set at 0.996.
Proposed modules: The effectiveness of semi-supervised learning depends largely on the quality of pseudo-labels. Nevertheless, it is clear that the previously mentioned procedure does not take into account the varying impact of individual samples on training. This paper concentrates on two critical aspects that are directly related to the generation of pseudo-labels: the initial pair of unlabeled images, and the efficiency of the pseudo-label-generation network, aka the teacher model, in identifying changes. To offer more reliable supervisory signals to unlabeled information and reduce training uncertainty, we propose an adaptive training strategy to tackle these two key issues. Our initial proposal is to develop a metric that quantifies the uncertainty of pseudo-labels, serving as the basis for adaptive modifications. Following that, we recommend applying adaptive modifications at the image level to the unlabeled training samples and suitably integrating reliable contents. Additionally, we apply adaptive and selective EMA updates to the teacher network during the training phase to minimize variations, ensuring more consistent and higher-quality pseudo-labels. The details of these methods are discussed in the following sections.